The formula s = 16t^2 is used to approximate the distance s, in feet, that an object falls freely (from rest) in t seconds. Use this formula to solve the problem. (Round answer to the nearest tenth.)
A branch of a tree hangs 575 ft above the soil. If a twig breaks off, how long will it take it, falling freely, to reach the soil?

Answer by mananth(16946)

s=16t^2
575 = 16t62
575 / 16 = T^2
SQRT(575/16) =t
t= 5.99 seconds